OpenStudy (anonymous):

Find (d^2y/dx^2) in terms of x and y. 1-x^2y=x-y

OpenStudy (anonymous):

@parkjc06 where r u getting confused at?

OpenStudy (anonymous):

is it just asking me for the derivative of both sides?

OpenStudy (anonymous):

you have to take the derivative in both sides it is a double derivative

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ d }{ dx } (1-x^2y) = \frac{ d }{ dx } (x-y)\] \[0-2*x*y-x^2\frac{ dy }{ dx } = 1 - \frac{ dy }{ dx }\]

OpenStudy (anonymous):

do u follow the steps?

OpenStudy (anonymous):

after this step u need to derivate again with respect to x

OpenStudy (anonymous):

so the second derivative would be 1-2x=0 right?

OpenStudy (anonymous):

no first of all u need to simplify the above expression\[\frac{ dy }{ dx } - x^2\frac{ dy }{ dx } = 1 +2*x*y\] \[\frac{ dy }{ dx }(1-x^2) = 1 +2*x*y\]

OpenStudy (anonymous):

now take the second derivative and apply product rule wherever necessary

OpenStudy (anonymous):

well to complete the first derivative it would by dy/dx=[1+2xy]/[1-x^2]

OpenStudy (anonymous):

wouldn't that be quotient rule to find second derivative

OpenStudy (anonymous):

so you would end up with y"=(1)(2*dy/dt)-(1+2xy)(-2X)/(1-x^2)?